Ernie Manufacturing has projected sales of $155.5 million next year. Costs are expected to be $100 million and net investment is expected to be $17.5 million. Each of these values is expected to grow at 14 percent the following year, with the growth rate declining by 2 percent per year until the growth rate reaches 6 percent where it will remain. There are 5.5 million shares of stock outstanding. Investors require a return of 13 percent and the corporate tax rate is 40 percent. What is your estimate of the current stock price?
1
Expert's answer
2019-09-27T09:05:16-0400

The formula for current stock price is:

P = D1/(r - g),

D1 = (155.5 - 100 - 17.5)×(1 - 0.4)×0.14 = 25.99.

P = 25.99/(0.13 - 0.12) = 2599.